OpenStudy (anonymous):

Can someone please help...Mars, Inc. claims that 24% of its M&M plain candies are blue. A sample of 100 M&Ms is randomly selected. Find the mean and standard deviations for the number of blue M&Ms in such a group of randomly selected 100 candies. This is what I have so far: Mean = np = 100(.24) = 24 Stdev = 100 x .24 x

OpenStudy (kropot72):

The standard deviation for a binomial distribution is found as follows: \[s.d.=\sqrt{np(1-p)}\]\[=\sqrt{100\times0.24(1-0.24)}=you\ can\ calculate\]

OpenStudy (kropot72):

@brettwlvn Do you understand?

OpenStudy (anonymous):

i took a minute to eat something...sorry. I didn't think I had gone far enough.

OpenStudy (anonymous):

Were my calculations correct?

OpenStudy (kropot72):

Your calculation for the expected mean in the random sample of 100 is correct.

OpenStudy (anonymous):

Ok, thanks. I'm going back to school at the age of 50; it's been some time since my last math class! =]

OpenStudy (kropot72):

You're welcome :) Good wishes!

OpenStudy (anonymous):

Might you have a moment to help me work this through to the end?

OpenStudy (anonymous):

is it... Stdev = 100 x .24 x .76 = 18.24 ?

OpenStudy (kropot72):

Not really. Did you notice the radical symbol, which means you need to take the square root of 18.24 to find the standard deviation. Note: 18.24 is the variance, which is not requested.